OpenStudy (anonymous):

Evaluate. Let n = 2. 8n + 3n - 7 ------------------- 6*2n - 3(n -1) Enter the answer in lowest terms.

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (daniellelovee):

substitution

OpenStudy (anonymous):

7 - 3*2 = 1 + 16 = 17?

OpenStudy (daniellelovee):

no..remember PEMDAS

OpenStudy (anonymous):

7-3 = 4 *2?

OpenStudy (daniellelovee):



OpenStudy (anonymous):

4*2 = 8 * 8 = 64 * 2 = 128

OpenStudy (daniellelovee):



OpenStudy (daniellelovee):

and the just substract by the greatest common factor of the fraction (3) to get 5/7

OpenStudy (anonymous):

Thanks I got 2 more

OpenStudy (daniellelovee):

ok just post them and tag me

OpenStudy (daniellelovee):

btw @polarpenguin thanks for the medal!

OpenStudy (anonymous):

OKay

OpenStudy (daniellelovee):

and thank you too for the medal lol